The equilibrium constant for the gaseous reaction
H2 + I2 <=> 2HI is 50.2 at 448°c
Calculate the number of grams of HI that are in equilibrium with 1.25 mole of H2 and 63.5g of iodine at this temperature
1
Expert's answer
2020-02-28T12:13:28-0500

The constant for the equilibriumH2+I2↔2HI is given by the following expression:

"K=(HI) ^2\/(H2) (I2)" =50.2.

The amount of moles of iodine is:"n(I2)=m(I2)\/M(I2)" = 63.5(g)/254 (g/mole) = 0.25 moles.

 Using expression for the equilibrium constant one can find the amount of moles of HI that are in the equilibrium with given amounts of H2 and I2:

 n^2(HI) / (125*0.25) =50.2.

50.2 =n^2(HI)/0.3215

N^2(HI)=50.2*0.3125=15.7

N(HI)=3.96mol.

 The mass of the HI can be calculated in the following way by multiplying the amount of moles by the molar mass:

"m(HI)=n(HI)\u00d7M(HI)" = 3.96 (mole) × 128 (g/mole) = 507 g.